OpenStudy (anonymous):

What is the means-to-MAD ratio of the two data sets, expressed as a decimal? ___ Data set Mean Mean absolute deviation (MAD) 1 4.3 1 2 4.9 1.2

OpenStudy (anonymous):

@iambatman @phi

OpenStudy (anonymous):

Sorry. Can't help you. I would if I could. I recently posted a question an received absolutely no help. Some people are utterly rude. Trust me, I would help you if I could.

OpenStudy (anonymous):

Yeah, I completely understand. :3

OpenStudy (anonymous):

The Data Set is 1 and 2, the mean is 4.3 and 4.9 and the mad is 1 and 1.2

OpenStudy (phi):

means-to-MAD ratio do you know how to write a ratio as a fraction? for example ratio a to b can be written as \(\frac{a}{b} \) can you write the means to MAD ratio?

OpenStudy (anonymous):

Uh.. I don't think so. Could you walk me through it?

OpenStudy (phi):

write the ratio as a fraction. first part of the ratio goes "up top", the second part of the ratio "goes below"

OpenStudy (anonymous):

So, it would be \[\frac{ 4 }{ 3 } and \frac{ 1 }{ 1.2 }?\]

OpenStudy (phi):

they want the mean to mad ratio what is the first part of that ratio?

OpenStudy (phi):

mean is the first part for your first data set find the mean. make that the top find the second part of the ratio (the mad) . make that the bottom what do you get ?

OpenStudy (anonymous):

So.. 4.3 and 1?

OpenStudy (phi):

yes, but write it as a fraction. first number (mean) goes up top, the other in the bottom

OpenStudy (anonymous):

\[\frac{ 4.3 }{ 1 }\]

OpenStudy (anonymous):

@phi

OpenStudy (phi):

yes, and that simplifies to 4.3 now do the same thing for data set 2, and simplify

OpenStudy (anonymous):

\[\frac{ 4.9 }{ 1.2 }\]

OpenStudy (phi):

the "line" means divide. for example \[ \frac{4}{2} = 2 \]

OpenStudy (anonymous):

4.08333333 is what I got dividing it.

OpenStudy (anonymous):

and 4.3

OpenStudy (anonymous):

So now what do I do?

OpenStudy (phi):

do you have answer choices?

OpenStudy (phi):

I have not seen this problem before. Do you have notes on this ?

OpenStudy (anonymous):

Sadly, no.

OpenStudy (phi):

I need a definition of "means-to-MAD" ratio of 2 data sets. I was assuming they wanted the two numbers you found 4.3 and 4.08 but if they want one number, we need to know what they are asking

OpenStudy (phi):

Your book must mention this. can you post a copy of the section that talks about this?